OpenStudy (anonymous):

Jernel has to figure out the area of her square garage. She knows that one side of the garage is equal to the length of her rabbit pen. The dimensions of the rectangular rabbit pen are 18 by 10.What is the area of the garage?

OpenStudy (pi123):

I can try to help but i cant gaurantee that it will be the right answer. ill try!

OpenStudy (anonymous):

ok ty

OpenStudy (pi123):

your welcome! give me a minute

OpenStudy (pi123):

a=l*w so... a=18*10 I hope this helped :)

OpenStudy (anonymous):

thank you

OpenStudy (pi123):

your welcome!